# Geometric Bounds for Persistence

### v1

@@ 2.1.4 Homological Spectra
As described in the introduction, the chief goal of our paper is to provide effective bounds for the lifetime of homology classes that appear along a (geometric) filtration of a metric space. We will formulate and realize this goal in a setting that encompasses, but is more general than, persistence diagrams.
We recall some additional definitions and results from [LMO23].
@@
$\mathrm{Spec}_{k}(\Delta_{\bullet}):=\bigcup_{r>0}\bigg{(}\operatorname{\tilde{H}}_{k}(\Delta_{r};\mathbb{F})\backslash\{0\}\times\{r\}\bigg{)}$ (1)
be the $k$-th homological spectrum of $\Delta_{\bullet}$ (with coefficients in $\mathbb{F})$.
@@
$b_{(\omega,s)}:=\inf\{r>0:r\leq s\text{ and }\exists\text{ nonzero }\omega_{r}\in\operatorname{\tilde{H}}_{k}(\Delta_{r})\text{ such that}(\iota_{r,s})_{\ast}(\omega_{r})=\omega\},$ (2)
$d_{(\omega,s)}:=\sup\{t>0:t\geq s\text{ and }\exists\text{ nonzero }\omega_{t}\in\operatorname{\tilde{H}}_{k}(\Delta_{t})\text{ such that}(\iota_{s,t})_{\ast}(\omega)=\omega_{t}\}.$ (3)
Whenever $\Delta_{\bullet}$ is a neighborhood filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$, as in [LMO23, Theorem 8], one has that $b_{(\omega,s)}<s\leq d_{(\omega,s)}$. [^6] In general, the type of intervals (open-open,closed-open, etc) one obtains depends on whether the filtration is defined via open or closed neighborhoods. Note that we’ve defined neighborhood filtrations via open neighborhoods. Let
@@
Informally, the interval $I_{(\omega,s)}$ encodes the maximal region around $s\in\mathbb{R}_{>0}$ inside which the class $\omega$ is “alive”.777 Note that, since we are using reduced homology, there is no degree 0 class $\omega$ such that $d_{(\omega,s)}=\infty$.
@@
Remark 2.19.
We focus on the special case of neighborhood filtrations. The birth time $b_{(\omega,s)}$ was defined as the infimum of all times $r\leq s$ when a “predecessor” of $\omega$ exists in $\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$, and it is natural to ask if there is a homology class that is supported on
@@
and also homologous to $\omega$ in $\mathcal{N}_{s}(\mathcal{X}\subset\mathbf{Y})$. It turns out this is not the case even for the neighborhood filtration of a compact set in $\mathbb{R}^{2}$. Namely, there exists a compact set $\mathcal{X}\subset\mathbb{R}^{2}$, known as the Warsaw circle or closed topologist’s sine curve [Bor75], satisfying the following counter-intuitive property: its first singular homology is zero, but every open neighborhood of it contains a homologically non-trivial circle. Every two of those circles are homologous to each other (within the union of the two neighborhoods). But these circles do not converge, as we shrink the neighborhood, to a non-trivial homology class of $\mathcal{X}$, because $\operatorname{\tilde{H}}_{1}(\mathcal{X})=0$. The natural way to treat this “limit circle” is to consider the Čech homology instead of the ordinary singular homology. The Čech homology $\mathrm{\check{H}}_{1}(\mathcal{X})$ is non-trivial, and contains the “limit circle”. In general, a predecessor of the class $\omega$ naturally lives in the Čech homology $\mathrm{\check{H}}_{k}(\overline{\mathcal{N}}_{b_{(\omega,s)}}(\mathcal{X}\subset\mathbf{Y}))$.
@@
to be closely related to the $k$-th persistence diagram of $\Delta_{\bullet}$. Whereas Proposition 2.21 below establishes a sense in which this is the case, it is not always true that all such intervals appear in the interval decomposition of $\operatorname{\tilde{H}}_{k}(\Delta_{\bullet})$ (whenever it exists). An example showing this discrepancy in the case of Vietoris-Rips filtrations can be found in [LMO23, Example 9.2].
Proposition 2.21 ([LMO23, Proposition 9.2]).
@@
This proposition indicates that for each interval $I=(r,s]$ in the barcode of $\Delta_{\bullet}$ there is a finite linearly independent collection $\omega_{1},\ldots,\omega_{m}\in\mathrm{H}_{k}(\Delta_{s})$ satisfying the conditions above such that $I=I_{(\omega_{i},s)}$ for all $i$. One calls any such $\omega_{i}$ a representative of the interval $I$ [^8] In [LMO23, Proposition 9.2] the authors consider the case of the VR filtration of a totally bounded metric space. The same proof applies to the more general statement given above. See [LMO23, page 42] for an example demonstrating the role of the condition that no nonzero linear combination of these vectors belongs to $\mathrm{Im}((\iota_{r,s})_{*})$.
@@
The above proposition implies that if $I$ is an interval in the barcode of a neighborhood filtration, then its length is equal to that of $I_{\omega_{i}}$, where $\omega_{i}$ is as in the statement. Therefore, and as we will do in the rest of the paper, if we have an upper bound for the length of all intervals $I_{\omega}$ where $\omega\in\mathrm{Spec}_{k}(\Delta_{\bullet})$ then we will automatically have an upper bound for the length of every interval in the barcode of $\Delta_{\bullet}$.
Throughout the paper, all homology coefficients are taken as $\mathbb{Z}_{2}$, unless otherwise stated. All results hold for any other coefficient field.
@@ 2.2 Geometry of $L^{\infty}(\mathcal{X})$ and Tight Spans
A metric space $E$ is called hyperconvex if for every family $(x_{i},r_{i})_{i\in I}$ of $x_{i}$ in $E$ and $r_{i}\geq 0$ such that $\mathbf{d}_{E}(x_{i},x_{j})\leq r_{i}+r_{j}$ for every $i,j$ in $I$, there exists a point $x\in E$ such that $\mathbf{d}_{E}(x_{i},x)\leq r_{i}$ for every $i$ in $I$.
This implies that the Čech and VR filtrations of a subset of $L^{\infty}(\mathcal{X})$ coincide (up to a factor of two in the filtration index). This gives a way to study VR-lifespans of homology of $\mathcal{X}$ by immersing it in $L^{\infty}(\mathcal{X})$ and then using the tools applicable for Čech-lifespans. Note that in the context of VR-persistence, this idea has been successfully used, cf. 2.32.
The hyperconvexity implies certain universal properties of $L^{\infty}$ spaces, and these properties will be implicitly used below in relation to Urysohn width (see 3.14), and tight spans (a.k.a. injective hulls). We briefly discuss the latter now (for more information we refer the reader to [Lan13]).
@@
A metric space $E$ is called injective if for every $1$-Lipschitz map $\phi:\mathcal{X}\to E$ and distance preserving embedding of $\mathcal{X}$ into $\widetilde{\mathcal{X}}$, there exists a $1$-Lipschitz map $\widetilde{\phi}:\widetilde{\mathcal{X}}\to E$ extending $\phi$:
[…]
@@
The proof of this proposition can be found in  [AP56] and [Lan13, Proposition 2.3].
@@
Tight spans are sometimes called injective hulls or hyperconvex hulls.
@@
    $\operatorname{E}(\mathcal{X})\cong\mathcal{X}$ for any metric tree $\mathcal{X}$.
@@
Note that from the characterization above, for any $f\in\operatorname{E}(\mathcal{X})$ and any $x\in\mathcal{X}$ it holds that
$f(x)=\max_{x\in\mathcal{X}}\big{(}\mathbf{d}_{\mathcal{X}}(x,x^{\prime})-f(x^{\prime})\big{)}=\big{\|}\mathbf{d}_{\mathcal{X}}(x,\cdot)-f\big{\|}_{\infty}.$ (4)
@@
Not only can the tight span $\operatorname{E}(\mathcal{X})$ be regarded as a subset of $L^{\infty}(\mathcal{X})$ but also, according to [Lan13, Proposition 2.2], there is a $1$-Lipschitz retraction of $L^{\infty}(\mathcal{X})$ to $\operatorname{E}(\mathcal{X})$.
@@ 2.3 Filling Radius
The filling radius is a key notion in metric geometry introduced by Gromov as a measure of largeness of a given closed manifold [Gro83]. To fill a manifold $\mathcal{M}$ of dimension $n$, treated as a singular $n$-cycle, means to find an $(n+1)$-dimensional singular chain $\mathcal{D}$ with boundary $\mathcal{M}$, i.e. $\partial\mathcal{D}=\mathcal{M}$. The coefficients of singular homology should be specified, as we discuss below.
@@
The filling radius $\rho(\mathcal{M})$ of a closed $n$-dimensional Riemannian manifold $\mathcal{M}$ is the infimal number $R>0$ such that $\mathcal{M}$ can be filled inside of the $R$-neighborhood $\mathcal{N}_{R}(\kappa(\mathcal{M}))$ of its Kuratowski image in $L^{\infty}(\mathcal{M})$.
In this definition, the most common convention is to use $\mathbb{Z}$-coefficients if $\mathcal{M}$ is oriented, and $\mathbb{Z}_{2}$-coefficients if $\mathcal{M}$ is not oriented. However, the usual persistence homology is well defined [^9] See, however, [Pat18]. only over fields, so we stick to $\mathbb{Z}_{2}$-coefficients. We note that all homology coefficients in the statements below are $\mathbb{Z}_{2}$, unless otherwise stated.
Before Gromov, a different type of filling radius notion was discussed for submanifolds (or more generally, cycles) in $\mathbb{R}^{N}$ in geometric analysis, especially in the context of isoperimetric problem [BS83, FF60, MS73]. We summarize both types of filling radii in the following definition, adapting it to the context of persistence.
Definition 2.31 (Relative and absolute filling radius of a homology class).
Let $(\mathcal{X},\mathbf{d}_{\mathcal{X}})$ be a metric space, and let $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$ be a non-trivial reduced singular homology class.
@@
    Assume additionally that $\mathcal{X}$ is a subset of a Banach space $\mathbf{Y}$, so that the metric $\mathbf{d}_{\mathcal{X}}$ agrees with the one inherited from $\mathbf{Y}$ (that is, the embedding $\mathcal{X}\hookrightarrow\mathbf{Y}$ is distance-preserving). The relative filling radius of $\omega$ is the infimal number $r$ such that the image of $\omega$ under the induced map $\operatorname{\tilde{H}}_{k}(\mathcal{X})\to\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}))$ is trivial. In other words, it is the infimal size of a neighborhood in $\mathbf{Y}$ where some cycle representing $\omega$ bounds a $(k+1)$-chain. Notation for relative filling radius: $\rho(\omega;\mathcal{X}\subset\mathbf{Y})$.
@@
    In case when no ambient space $\mathbf{Y}$ is specified, one can take $\mathbf{Y}=L^{\infty}(\mathcal{X})$, and measure the relative filling radius of the Kuratowski image $\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})$. This way one gets the (absolute) filling radius of $\omega$: $\rho(\omega;\mathcal{X})=\rho(\kappa_{*}\omega;\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$; cf. [LMO23, Definition 24].
The word “absolute”, which is usually omitted, is justified by the fact that $\rho(\omega;\mathcal{X})$ equals the minimum of $\rho(\iota_{*}\omega;\iota(\mathcal{X})\subset\mathbf{Y})$ over all distance-preserving embeddings $\iota:\mathcal{X}\hookrightarrow\mathbf{Y}$ [Gro83]; we will refer to this as the universal property. Note that when $\mathcal{X}=\mathcal{M}$ is a closed Riemannian manifold, Gromov’s filling radius (2.30) over $\mathbb{Z}_{2}$ is the same as the absolute filling radius of the $\mathbb{Z}_{2}$-fundamental class $[\mathcal{M}]$. Note also that the relationship between the VR filtration of a manifold and its Gromov’s filling radius is studied in [LMO23, Section 9.3].
Proposition 2.32 ([LMO23, Proposition 9.4]).
Let $\mathcal{M}$ be an $n$-dimensional closed connected Riemannian manifold. Then, the interval $I_{\mathcal{M}}:=(0,2\rho(\mathcal{M})]$ appears in the degree-$n$ barcode of $\mathcal{V}_{\bullet}(\mathcal{M})$. Here both the filling radius and persistent homology are computed with $\mathbb{Z}_{2}$ coefficients.
@@
    One way is to consider the intrinsic Riemannian metric $g$ on $\mathrm{E}$ induced by the Euclidean metric of $\mathbb{R}^{3}$. Thus obtained Riemannian surface $\mathrm{E}_{g}$ has the absolute filling radius about $10$: $\rho([\mathrm{E}];\mathrm{E}_{g})\sim 10$ (the exact computation is tricky).
@@
In the following sections, we will see that Gromov’s filling radius of a Riemannian manifold isometrically embedded in an ambient space $\mathbf{Y}$ can highly overestimate the lifespan of a topological feature in the Čech filtration and that the relative filling radius is better adapted to this context.
Filling Radius & Persistence. The notion of lifespan of a homology class (2.18) is directly related to the filling radius as follows. Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Let $(b_{\omega},d_{\omega})$ be the homological birth and death times of a degree-$k$ homology class $\omega$ present at time $s$ in the neighborhood filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$. Right from the definition of the relative filling radius, we get
@@
A bit more generally, if $\omega_{r}\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ is a predecessor of $\omega$ (that is, it is mapped to $\omega$ by the map in homology induced by the inclusion), then
@@
Letting $r\to b_{\omega}$, we get a formula for the lifespan of $\omega$ in terms of the filling radii of the predecessors of $\omega$:
@@
As clarified in 2.19, this lifespan cannot be written in term of the filling radius of homology of $\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\mathbf{Y})$. We cannot expect (a predecessor of) $\omega$ to be present in the homology of $\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\mathbf{Y})$, unless we work with Čech homology, and redefine the filling radius correspondingly.
@@
    $d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})).$

### v2

@@ 2.1.4 Homological Spectra
As described in the introduction, the chief goal of our paper is to provide effective bounds for the lifetime of all homology classes that appear along a (geometric) filtration of a metric space. We will formulate and realize this goal in a setting that encompasses, but is more general than, persistence diagrams.
Remark 2.17.
Notice that it is not true that the only homology classes that show up across the filtration are those coming from the initial space. One well known example is that of the circle $S^{1}$ (with its geodesic distance) and the Vietoris–Rips filtration. Indeed, as shown by Adamaszek and Adams in [AA17], whereas $\mathcal{V}_{t}(S^{1})$ has the homotopy type of $S^{1}$ for $t\in(0,\tfrac{2\pi}{3}]$, the homotopy type is that of $S^{3}$ as soon as $t\in(\tfrac{2\pi}{3},\tfrac{4\pi}{5}]$. In fact, they show that $\mathcal{V}_{\bullet}(S^{1})$ eventually attains the homotopy types of all odd-dimensional spheres.
We recall some additional definitions and results from [LMO24].
@@
$\mathrm{Spec}_{k}(\Delta_{\bullet}):=\bigcup_{r>0}\bigg{(}\widetilde{\mathrm{H}}_{k}(\Delta_{r};\mathbb{F})\backslash\{0\}\times\{r\}\bigg{)}$ (1)
[…]
@@
$b_{(\omega,s)}$ $:=\inf\{r>0:r\leq s\text{ and }\exists\text{ nonzero }\omega_{r}\in\widetilde{\mathrm{H}}_{k}(\Delta_{r})\text{ such that }(\iota_{r,s})_{\ast}(\omega_{r})=\omega\},$ (2)
$d_{(\omega,s)}$ $:=\sup\{t>0:t\geq s\text{ and }\exists\text{ nonzero }\omega_{t}\in\widetilde{\mathrm{H}}_{k}(\Delta_{t})\text{ such that }(\iota_{s,t})_{\ast}(\omega)=\omega_{t}\}$ (3)
$\,\,=\sup\{t\geq s:(\iota_{s,t})_{\ast}(\omega)\neq 0\}.$ (4)
Whenever $\Delta_{\bullet}$ is a neighborhood filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$, as in [LMO24, Theorem 8], one has that $b_{(\omega,s)}<s\leq d_{(\omega,s)}$. [^7] In general, the type of intervals (open-open, closed-open, etc) one obtains depends on whether the filtration is defined via open or closed neighborhoods. Note that we’ve defined neighborhood filtrations via open neighborhoods. Let
@@
Informally, the interval $I_{(\omega,s)}$ encodes the maximal region around $s\in\mathbb{R}_{>0}$ inside which the class $\omega$ is “alive”.888 Note that, since we are using reduced homology, there is no degree 0 class $\omega$ such that $d_{(\omega,s)}=\infty$.
@@
Remark 2.20 (A Caveat).
We focus on the special case of neighborhood filtrations. The birth time $b_{(\omega,s)}$ was defined as the infimum of all times $r\leq s$ when a “predecessor” of $\omega$ exists in $\widetilde{\mathrm{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$, and it is natural to ask whether there exists a homology class supported on
@@
that is also homologous to $\omega$ in $\mathcal{N}_{s}(\mathcal{X}\subset\mathbf{Y})$. It turns out this is not the case even for the neighborhood filtration of a compact set in $\mathbb{R}^{2}$. Namely, there exists a compact set $\mathcal{X}\subset\mathbb{R}^{2}$, known as the Warsaw circle or closed topologist’s sine curve, satisfying the following counter-intuitive property: its first singular homology is zero, but every open neighborhood of it contains a homologically non-trivial circle; see Borsuk [Bor75]. Every two of those circles are homologous to each other (within the union of the two neighborhoods). But these circles do not converge, as we shrink the neighborhood, to a non-trivial homology class of $\mathcal{X}$, because $\widetilde{\mathrm{H}}_{1}(\mathcal{X})=0$. The natural way to treat this “limit circle” is to consider Čech homology instead of ordinary singular homology. The Čech homology $\widecheck{\mathrm{H}}_{1}(\mathcal{X})$ is non-trivial, and contains the “limit circle”. In general, a predecessor of the class $\omega$ naturally lives in the Čech homology $\widecheck{\mathrm{H}}_{k}(\overline{\mathcal{N}}_{b_{(\omega,s)}}(\mathcal{X}\subset\mathbf{Y}))$. Nonetheless, we consistently use singular homology throughout the paper in order to avoid overly technical details.
@@
[…]
Proposition 2.22 ([LMO20, Proposition 9.2]).
@@
This proposition indicates that for each interval $I=(r,s]$ in the barcode of $\Delta_{\bullet}$ there is a finite linearly independent collection $\omega_{1},\ldots,\omega_{m}\in\mathrm{H}_{k}(\Delta_{s})$ satisfying the conditions above such that $I=I_{(\omega_{i},s)}$ for all $i$. One calls any such $\omega_{i}$ a representative of the interval $I$ [^9] In [LMO24, Proposition 9.2] the authors consider the case of the VR filtration of a totally bounded metric space. The same proof applies to the more general statement given above. See [LMO24, page 42] for an example demonstrating the role of the condition that no nonzero linear combination of these vectors belongs to $\mathrm{Im}((\iota_{r,s})_{*})$.
@@
The above proposition implies that if $I$ is an interval in the barcode of a neighborhood filtration, then $I=I_{\omega_{i}}$ so that, in particular, its length is equal to that of $I_{\omega_{i}}$, where $\omega_{i}$ is as in the statement. Therefore, and as we will do in the rest of the paper, if we have an upper bound for the length of all intervals $I_{\omega}$ where $\omega\in\mathrm{Spec}_{k}(\Delta_{\bullet})$ then we will automatically have an upper bound for the length of every interval in the barcode of $\Delta_{\bullet}$.
@@ 2.2 Geometry of $L^{\infty}(\mathcal{X})$ and Tight Spans
A metric space $(\mathcal{E},\mathbf{d}_{\mathcal{E}})$ is called hyperconvex if for every family $(x_{i},r_{i})_{i\in I}$ of $x_{i}$ in $\mathcal{E}$ and $r_{i}\geq 0$ such that $\mathbf{d}_{\mathcal{E}}(x_{i},x_{j})\leq r_{i}+r_{j}$ for every $i,j$ in $I$, there exists a point $x\in\mathcal{E}$ such that $\mathbf{d}_{\mathcal{E}}(x_{i},x)\leq r_{i}$ for every $i$ in $I$.
This implies that the Čech and VR filtrations of a subset of $L^{\infty}(\mathcal{X})$ coincide (up to a factor of two in the filtration index). This gives a way to study VR-lifespans of homology of $\mathcal{X}$ by immersing it in $L^{\infty}(\mathcal{X})$ and then using the tools applicable for Čech-lifespans. Note that in the context of VR-persistence, this idea has been successfully used, cf. 2.34.
The hyperconvexity implies certain universal properties of $L^{\infty}$ spaces, and these properties will be implicitly used below in relation to Urysohn width (see 3.15), and tight spans (a.k.a. injective hulls). We briefly discuss the latter now and we refer the reader to Lang’s survey [Lan13] for more information.
@@
A metric space $\mathcal{E}$ is called injective if for every $1$-Lipschitz map $\phi:\mathcal{X}\to\mathcal{E}$ and distance preserving embedding of $\mathcal{X}$ into $\widetilde{\mathcal{X}}$, there exists a $1$-Lipschitz map $\widetilde{\phi}:\widetilde{\mathcal{X}}\to\mathcal{E}$ extending $\phi$:
${\mathcal{X}}$${\widetilde{\mathcal{X}}}$${\mathcal{E}}$$\scriptstyle{\phi}$$\scriptstyle{\widetilde{\phi}}$
@@
The proof of this proposition can be found in Aronszajn and Panitchpakdi [AP56] and [Lan13, Proposition 2.3].
@@
Tight spans are sometimes called injective hulls or hyperconvex hulls. The notions of injectivity and hyperconvexity were first proposed by Aronszajn and Panitchpakdi [AP56]. Isbell [Isb64] first identified the notion of tight span (although the author used the term injective envelope). Additional contributions were made by Dress [Dre84] and Lang [Lan13]; see Chepoi [Che97] for a historical account.
@@
    $\operatorname{E}(\mathcal{X})$ is isometric to $\mathcal{X}$ for any metric tree $\mathcal{X}$.
@@
Note that from the realization of the tight span recalled above, for any $f\in\operatorname{E}(\mathcal{X})$ and any $x\in\mathcal{X}$, it holds that
$f(x)=\max_{x^{\prime}\in\mathcal{X}}\big{(}\mathbf{d}_{\mathcal{X}}(x,x^{\prime})-f(x^{\prime})\big{)}=\big{\|}\mathbf{d}_{\mathcal{X}}(x,\cdot)-f\big{\|}_{\infty}.$ (5)
@@
Not only can the tight span $\operatorname{E}(\mathcal{X})$ be regarded as a subset of $L^{\infty}(\mathcal{X})$ but also, directly from the fact that it is an injective metric space, there is a $1$-Lipschitz retraction of $L^{\infty}(\mathcal{X})$ to $\operatorname{E}(\mathcal{X})$; see [Lan13, Proposition 2.2].
@@ 2.3 Filling Radius
The filling radius is a key notion in metric geometry introduced by Gromov as a measure of largeness of a given closed manifold [Gro83]. To fill a manifold $\mathcal{M}$ of dimension $n$, treated as a singular $n$-cycle, means to find an $(n+1)$-dimensional singular chain $\mathcal{D}$ with boundary $\mathcal{M}$, i.e. $\partial\mathcal{D}=\mathcal{M}$. In this case we will also be saying that the cycle $\mathcal{M}$ bounds. The ambient space in which filling happens, as well as the coefficients of singular homology, should be specified, as we discuss below.
@@
The filling radius $\rho(\mathcal{M})$ of a closed $n$-dimensional Riemannian manifold $\mathcal{M}$ is the infimal number $R>0$ such that $\mathcal{M}$ can be filled inside of the $R$-neighborhood of its Kuratowski image in $L^{\infty}(\mathcal{M})$.
Remark 2.32 (Coefficients).
This definition makes sense with any homology coefficients. Some common choices include $\mathbb{Z}$-coefficients if $\mathcal{M}$ is oriented, $\mathbb{Z}_{2}$-coefficients if $\mathcal{M}$ is not oriented, and $\mathbb{Q}$-coefficients in some contexts where torsion is a problem. However, the usual persistence homology is well-defined only over fields, [^10] See, however, [Pat18]. so in the rest of the paper we implicitly assume that an arbitrary choice of a field is made (e.g., $\mathbb{Z}_{2}$), and that all filling radii and all persistence features are considered over this field. All of our results hold for any field and, for this reason, the field is omitted from the notation.
Before Gromov, a different type of filling radius notion was discussed for submanifolds (or more generally, cycles) in $\mathbb{R}^{N}$ in geometric analysis, especially in the context of the isoperimetric problem [FF60, MS73, BS83]. We summarize both types of filling radii in the following definition, adapting it to the context of persistence.
Definition 2.33 (Relative and Absolute Filling Radii of a Homology Class).
Let $(\mathcal{X},\mathbf{d}_{\mathcal{X}})$ be a metric space, and let $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{X})$ be a non-trivial reduced singular homology class.
@@
    Assume additionally that $\mathcal{X}$ is a subset of a Banach space $\mathbf{Y}$, so that the metric $\mathbf{d}_{\mathcal{X}}$ agrees with the one inherited from $\mathbf{Y}$ (that is, the embedding $\mathcal{X}\hookrightarrow\mathbf{Y}$ is distance-preserving). The relative filling radius of $\omega$ is the infimal number $r$ such that the image of $\omega$ under the induced map $\widetilde{\mathrm{H}}_{k}(\mathcal{X})\to\widetilde{\mathrm{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ is trivial. In other words, it is the infimal size of a neighborhood in $\mathbf{Y}$ where some cycle representing $\omega$ bounds a $(k+1)$-chain. Notation for relative filling radius: $\rho(\omega;\mathcal{X}\subset\mathbf{Y})$.
@@
    In case when no ambient space $\mathbf{Y}$ is specified, one can take $\mathbf{Y}=L^{\infty}(\mathcal{X})$, and measure the relative filling radius of the Kuratowski image $\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})$. This way one obtains the (absolute) filling radius of $\omega$: $\rho(\omega;\mathcal{X})=\rho(\kappa_{*}(\omega);\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$; cf. [LMO24, Definition 24].
The word “absolute”, which is usually omitted, is justified by the fact that $\rho(\omega;\mathcal{X})$ equals the infimum of $\rho(\iota_{*}(\omega);\iota(\mathcal{X})\subset\mathbf{Y})$ over all distance-preserving embeddings $\iota:\mathcal{X}\hookrightarrow\mathbf{Y}$ [Gro83, page 8]; we will refer to this as the universal property. Note that when $\mathcal{X}=\mathcal{M}$ is a closed Riemannian manifold, Gromov’s filling radius (2.31) over $\mathbb{Z}_{2}$ is the same as the absolute filling radius of the $\mathbb{Z}_{2}$-fundamental class $[\mathcal{M}]$.
The relationship between the VR filtration of a compact metric space and the absolute filling radii $\rho(\omega;\mathcal{X})$ is studied in [LMO24, Section 9.3]. The following proposition applies in the general context of absolute neighborhood retracts (ANRs), which includes Riemmanian manifolds, metric graphs and other commonly appearing metric spaces. [^11] Recall that an ANR is any metric space $\mathcal{X}$ with the property that whenever it is embedded into another metric space $\mathcal{Z}$ through a homeomorphism $h:\mathcal{X}\to\mathcal{Z}$, then there is an open neighborhood $U$ of $h(\mathcal{X})$ such that $h(\mathcal{X})$ is a retract of $U$; see Borsuk [Bor32] and Hu [Hu65]. It is known that every compact (topologically) finite-dimensional locally contractible metric space is an ANR. Thus, all Riemannian manifolds are ANRs.
Proposition 2.34 ([LMO24, Propositions 9.28 and 9.46]).
Let $\mathcal{X}$ be a compact ANR metric space. Then, for any integer $k\geq 1$ and any nonzero $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{X})$ we have:
  • $\rho(\omega;\mathcal{X})>0$;
  • the interval $(0,2\rho(\omega;\mathcal{X})]$ appears in the degree-$k$ barcode of $\mathcal{V}_{\bullet}(\mathcal{X}).$
Additionally, if $\mathcal{X}$ is a (closed and connected) Riemannian manifold, then there are no other intervals with left-endpoint (birth time) equal to zero in the degree-$k$ barcode of $\mathcal{V}_{\bullet}(\mathcal{X})$. [^12] In particular, if $\mathcal{X}$ is an $n$-dimensional Riemannian manifold, then $(0,2\rho(\mathcal{X})]$ is the unique interval with left endpoint zero in the degree-$n$ barcode. Here both the filling radius and persistent homology can be computed with coefficients in an arbitrary field when $\mathcal{M}$ is orientable, and with coefficients in $\mathbb{Z}_{2}$ when $\mathcal{M}$ is not orientable.
@@
    One way is to consider the Riemannian metric $g$ on $\mathrm{E}$ induced by the Euclidean metric of $\mathbb{R}^{3}$. Thus obtained Riemannian surface $\mathrm{E}_{g}$ has the absolute filling radius of about $10$: $\rho([\mathrm{E}];\mathrm{E}_{g})\sim 10$ (the exact computation is tricky).
@@
In the following sections, we will see that Gromov’s filling radius of a Riemannian manifold isometrically embedded in an ambient space $\mathbf{Y}$ can highly overestimate the lifespan of a topological feature in its Čech filtration and that the relative filling radius is better adapted to this context.
Remark 2.36 (Comment about Estimation of Filling Invariants).
The type connections between persistent homology and metric geometry that we explore in this paper have the potential of permitting the estimation of quantities such as the filling radii $\rho(\omega;\mathcal{X})$, $\omega\in\widetilde{\mathrm{H}}_{k}(\mathcal{X})$, via the polynomial time algorithms that have been developed for computing persistent homology (see Section 2.1). Indeed, such an estimate would be obtained via 2.34 and the stability of PH (2.15) through computing the VR-barcodes of a carefully chosen $\epsilon$-net for $\mathcal{X}$, for some $\epsilon>0$.
Filling Radii & Lifespans. The notion of lifespan of a homology class (2.19) is directly related to the filling radius as follows. Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. Let $(b_{\omega},d_{\omega})$ be the homological birth and death times of a degree-$k$ homology class $\omega$ present at time $s$ in the neighborhood filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$. Directly from the definition of the relative filling radius, we obtain
@@
A bit more generally, if $\omega_{r}\in\widetilde{\mathrm{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ is a predecessor of $\omega$ (that is, it is mapped to $\omega$ by the map in homology induced by the inclusion), then
@@
Letting $r\to b_{\omega}$, we obtain a formula for the lifespan of $\omega$ in terms of the filling radii of the predecessors of $\omega$:
@@
Note that both sides of this formula depend on homology coefficients lying in the same field, which can be arbitrary.
Remark 2.37.
As clarified in 2.20, this lifespan cannot be written in term of the filling radius of homology of $\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\mathbf{Y})$. We cannot expect (a predecessor of) $\omega$ to be present in the homology of $\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\mathbf{Y})$, unless we work with Čech homology, and redefine the filling radius correspondingly.
@@
    $d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho\big{(}\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})\big{)}.$